OpenStudy (sierra!):

A tent has the shape of a triangular prism. Each end of the tent is a triangle with base length 8 ft and height 5 ft. The tent is 12 ft long from front to back. What is the volume of the tent? A. 25 ft3 B. 240 ft3 C. 328 ft3 D. 480 ft3

OpenStudy (sierra!):

@Unanimoose

OpenStudy (unanimoose):

Ok, the same as before. The equation is for a traingular prism. \[V = \frac{ a }{ 2 } \times 8 \times 5\]

OpenStudy (unanimoose):

Where a = 12, c = 8 and h = 5.

OpenStudy (unanimoose):

Woops, I meant to put c and h but I guess I did that for you, hehe.

OpenStudy (sierra!):

so how do i do that

OpenStudy (unanimoose):

Replace a with 12, divide them and then multiply the values.

OpenStudy (sierra!):

so D

OpenStudy (sierra!):

i have one more this looks confusing

OpenStudy (sierra!):

What is the volume of the triangular prism? A. 7168 ft3 B. 3584 ft3 C. 1792 ft2 D. 62 ft3

OpenStudy (unanimoose):

Not quite, you forgot to divide 12 and 2~

OpenStudy (unanimoose):

For the other one tho, you didn't do 12/2. You came out with double what you wanted.

OpenStudy (unanimoose):

12/2 = 6 6*5*8 = 240

OpenStudy (unanimoose):

But anyway, attach the image to for this one. ^^

OpenStudy (sierra!):

so 240

OpenStudy (unanimoose):

Yeah.

OpenStudy (unanimoose):

\[v = \frac{ a }{ 2 } \times c \times h\] Try this one on your own. :)

OpenStudy (unanimoose):

I'll be attending another thread, I'll come back when you reply. ^^

OpenStudy (sierra!):

so 16/2=8

OpenStudy (sierra!):

8*14*32=3584

OpenStudy (sierra!):

so the answer is B

OpenStudy (sierra!):

Am i right @Unanimoose

OpenStudy (unanimoose):

Yep! Nice. ^^

OpenStudy (unanimoose):

All of them have been B so far, lol.

OpenStudy (sierra!):

THANK YOU I GOT 100

OpenStudy (sierra!):

can you fan me so we can text

OpenStudy (sierra!):

@Unanimoose

OpenStudy (unanimoose):

Os is a bit laggy atm so I can't fan anyone, but once it's calmed down a little I'll fan you. :P And no problem. :)